OpenStudy (anonymous):

Which of the following tables shows the correct steps to transform x2 + 8x + 15 = 0 into the form (x - p)2 = q? [p and q are integers]

OpenStudy (anonymous):

Answer choices: Step 1 x2 + 8x + 15 - 1 = 0 - 1 Step 2 x2 + 8x + 14 = -1 Step 3 (x + 4)2 = -1 Step 1 x2 + 8x + 15 - 2 = 0 - 2 Step 2 x2 + 8x + 13 = -2 Step 3 (x + 4)2 = -2 Step 1 x2 + 8x + 15 + 1 = 0 + 1 Step 2 x2 + 8x + 16 = 1 Step 3 (x + 4)2 = 1 Step 1 x2 + 8x + 15 + 2 = 0 + 2 Step 2 x2 + 8x + 17 = 2 Step 3 (x + 4)2 = 2

OpenStudy (anonymous):

@Squirrels

OpenStudy (squirrels):

I assume this x2 + 8x + 15 = 0 is this \[x^2+8x+15=0\]

OpenStudy (anonymous):

yeah

OpenStudy (triciaal):

no 3

OpenStudy (anonymous):

looks like you have the answer

OpenStudy (anonymous):

Oh I didn't even see that, yeah you do.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

the second one